OpenStudy (anonymous):

x^2 + 8x + 15 / x^2 - 3x - 18

OpenStudy (anonymous):

\[\frac{(x+3)(x+5)}{(x-6)(x+3)}\] \[= \frac{(x+5)}{(x-6)}\] Yes?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yup